# Efficient detection of entanglement by stimulated disentanglement

### v2

@@ 3 Discussion
The observed SNR in our measurements (37dB) underscores the critical role of time-energy entanglement in the SFG detection. The inherent coherence of the time-energy entangled photon pairs ensures that the measured SFG spectrum surpasses the shot-noise limit of the pump laser. Specifically, the noise of the pump field amplitude corresponds to one photon in amplitude ($1\sqrt{\text{photons}}/\text{mode}$), whereas the average field amplitude per mode in the experiment was only $\sim\!0.03\sqrt{\text{photons}}/\text{mode}$. Thus, the SFG amplitude of individual mode pairs was undetectable, substantially below this shot-noise limit, and yet, the coherent summation of all the modes across the SPDC spectrum generated a well detected SFG signal, >37dB above the noise floor.
This highlights that Coherent SFG detection acts as an ideal quantum matched filter that leverages the ultrafast coincidence nature of SFG and the energy-sum resolution to ideally reject accidental coincidences in both energy and time, dramatically enhancing the SNR and the detection sensitivity.
Thus, coherent SFG detection offers a fundamental enhancement over traditional detection schemes, such as SU(1,1) interferometry or coincidence / HOM, since SFG detection is global, combining contributions from the entire SPDC spectrum into a single SFG mode, whereas standard methods are local, independently measuring specific modes.
This global approach of coherent SFG and the dramatic SNR enhancement can be transformative for quantum sensing applications and broadband quantum technology, where the coherent summation opens a door to overcome major noise limitations. A major example is in applications of quantum illumination [39, 37], which exploit entanglement to enhance the detection of minute signals buried in classical noise, and offer substantial advantages for quantum radar systems, secure quantum communication, and advanced imaging techniques. Efficient SFG detection is ideal for realizing the practical benefits of quantum illumination in real-world applications.
@@ 4.1 Theoretical Model
where $\hat{N}_{s}=\sum_{\nu}\hat{a}_{\nu}^{\dagger}\hat{a}_{\nu}$ is the photon number over all the signal modes (and for the idler $\hat{N}_{i}$ at $-\nu$), $M$ is the number of interacting mode pairs and $\hat{H}^{\prime}=\hat{H}/\chi=\sum_{\nu}(\hat{a}_{p}^{\dagger}\hat{a}_{\nu}\hat{a}_{-\nu}+\hat{a}_{p}\hat{a}_{\nu}^{\dagger}\hat{a}_{-\nu}^{\dagger})$ is the unitless Hamiltonian. Equation 4 is a central result of the SFG analysis.
Let us pause briefly the derivation to consider the implications of Eq. 4: First, within the second order correction, SFG is only due to recombination of twin-modes back to the original pump frequency, whereas “accidental” (classical) SFG from uncorrelated modes will appear only in higher order corrections (4th and on). This is an indication of the quantum advantage of the twin-modes SFG over the classical picture. Second, the number of participating mode pairs $M$ reflects the spectral bandwidth of the PDC light $\Delta\nu$. In an OPO cavity, $M$ is simply the number of cavity comb-teeth within the gain bandwidth, dictated by phase matching. In our experiment, where the spectrum is continuous, $M\!=\!\Delta\nu T/2$ reflects the time-bandwidth product, where $\Delta\nu\!\approx\!113$THz and the effective measurement time $T\!=\!1/\delta\nu$ reflects the spectral resolution of our prism-based spectrometer $\delta\nu\!\approx\!0.5$THz, which indicates that the effective number of modes in our experiment was $M\!\approx\!60$. Clearly, the measurement time $T$ can be pushed much longer with a better spectrometer, ultimately limited only by the coherence time (linewidth) of the pump laser, which will push the number of modes to $M\!\approx\!10^{5-6}$ and beyond. And last, note that the only assumption thus far was that the signal-idler fields are weak relative to the pump, but the parametric gain was not limited, allowing $g\geq 1$, which indicates that Eq. 4 is applicable also to the high-squeezing multi-photon regime.
@@
[…]
@@
${\Delta N}_{p}=\frac{\mathrm{i}}{\sqrt{2}}{(\chi t)}^{3}\sum_{\nu}{\left[\alpha_{r}\hat{a}_{\nu}^{(in)\dagger}\hat{a}_{-\nu}^{(in)\dagger}-\alpha_{r}^{*}\hat{a}_{\nu}^{(in)}\hat{a}_{-\nu}^{(in)}\right]=\sqrt{2}\text{Im}\left[\alpha_{r}^{*}\hat{a}_{\nu}^{(in)}\hat{a}_{-\nu}^{(in)}\right]},$ (9)
@@ 4.2 Experimental Methodology
To generate an appreciable SFG signal from the entire SPDC bandwidth, it is crucial to maintain a flat spectral phase for all signal-idler frequency pairs in the ultra-broadband spectrum to ensure that every photon pair contributes constructively to the SFG process. This demands precise control of the group delay / dispersion in our SU(1,1) interferometer, achieved by a prism-based shaper of the spectral-phase in the PDC path. This phase shaper, which consists of a prism pair and a reflective 4f telescope, enables independent control of both the second and the fourth orders of dispersion (following the concept of [44]). In order to tune the dispersion to exact compensation, we measure the spectral phase in real time by observing the spectrum of the ultra-broadband SPDC at the output of the SU(1,1) interferometer with a home-built, prism-based spectrometer. The contrast of the observed SU(1,1) interference was optimized following the guidelines of [40].
@@
Finally, to isolate the desired SFG signal from the much stronger pump residue, we employ lock-in detection by dithering (modulating) the phase of the entire PDC spectrum (spontaneous and stimulated) according to $\phi(t)\!=\!\phi_{D}\cos\nu t$ using the PZT mirror in the spectral shaper (at frequency $\nu_{D}\!=\!2\pi\!\times\!173kHz$). Due to the coherent nature of the SFG process, this phase dithering is imprinted onto the SFG result, which then interferes with the pump-residue field, modulating the detected SFG intensity, as illustrated in Fig. 5c and analyzed theoretically later on. By measuring the spectral intensity of the detected SFG signal at the modulation frequency $I(\nu)$, we effectively isolate the modulated SFG contribution from the large and noisy DC background of the pump-residue.
@@ 4.3 Theoretical Calculation of Seeded Coherent SFG
[…]
@@
The experiment measured the RF spectrum of the SFG intensity, which is proportional to $\left\langle\Delta N_{p}\right\rangle^{2}$
@@ 5.1 The non-classical response to loss
To gain deeper insight into the detection dynamics of our coherent SFG method and the relations between the interfering SFG contributions - spontaneous and stimulated, we investigated the response of the SFG signal to SPDC loss, regarding the theoretical prediction of Eq. 12. SPDC loss affects both photons of a pair, leading to a unique quantum situation, where although the flux of individual photons is reduced by $T$ the biphoton rate is reduced quadratically, as the probability of both photons surviving the SPDC loss is $T^{2}$.
@@
Funding This project was funded by the Israel Innovation Authority and Elta Systems Ltd. under project No. 70002 and under a Danish-Israeli Eureka project. Funding is acknowledged also from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101178170 and from the Israel Science Foundation under grant No. 2208/24.

### v3

@@ 4 Discussion
By combining theory and experiment, we demonstrate that coherent SFG can act as an efficient quantum matched filter, providing a global joint readout of time-difference and frequency-sum correlations. The critical role of time-energy entanglement in the SFG detection is underscored by the fact that our measurements observed only the coherent SFG peak at the pump frequency with no background SFG at other frequencies (down to the dark noise of the RF spectrum analyzer). The inherent coherence of the time-energy entangled photon pairs ensures that the measured SFG spectrum surpasses the shot-noise limit of the pump laser. Specifically, the noise of the pump field amplitude corresponds to one photon in amplitude ($1\sqrt{\text{photons}}/\text{mode}$), whereas the average field amplitude per mode in the experiment was only $\sim\!0.03\sqrt{\text{photons}}/\text{mode}$. Thus, the SFG amplitude of individual mode pairs was undetectable, substantially below this shot-noise limit. And yet, the coherent summation of all the modes across the SPDC spectrum generated a well detected SFG signal, >37dB above the electronic noise floor.
This highlights that Coherent SFG detection acts as an ideal quantum matched filter which leverages the ultrafast coincidence nature of SFG and the energy-sum resolution to reject accidental coincidences efficiently in both energy and time, dramatically enhancing the SNR and the detection sensitivity.
Thus, coherent SFG detection offers a fundamental enhancement over traditional detection schemes, such as SU(1,1) interferometry or coincidence / HOM, since SFG detection is global, combining contributions from the entire SPDC spectrum into a single SFG mode, whereas standard methods are local, independently measuring specific modes. Our method demonstrates that entanglement detection is most effective through global measurement that probes the entire non-separable system, rather than by isolating individual components.
This approach of coherent SFG and the dramatic SNR enhancement can be transformative for quantum sensing applications and additional broadband quantum technologies, where the coherent summation opens a door to overcome significant noise limitations. A major example is quantum illumination [39, 37], which exploits continuous-variables entanglement to enhance the detection of minute signals buried in classical noise, thereby offering substantial advantages for quantum radar systems, secure quantum communication, and advanced imaging techniques. Efficient SFG detection is ideal for realizing the practical benefits of broadband entangled photons in general, and quantum illumination in particular, within noisy and lossy environments.
@@ 5.1 Theoretical Model
where $\hat{N}_{s}=\sum_{\nu}\hat{a}_{\nu}^{\dagger}\hat{a}_{\nu}$ is the photon number over all the signal modes (and for the idler $\hat{N}_{i}$ at $\nu$), $M$ is the number of interacting mode pairs and $\hat{H}^{\prime}=\hat{H}/\chi=\sum_{\nu}(\hat{a}_{p}^{\dagger}\hat{a}_{\nu}\hat{a}_{-\nu}+\hat{a}_{p}\hat{a}_{\nu}^{\dagger}\hat{a}_{-\nu}^{\dagger})$ is the unitless Hamiltonian. Eq. 7 is a central result of the SFG analysis.
Let us pause briefly the derivation to consider the implications of Eq. 7: First, within the second order correction, SFG is only due to recombination of twin-modes back to the original pump frequency, whereas “accidental” (classical) SFG from uncorrelated modes will appear only in higher order corrections (4th and on). This is an indication of the quantum advantage of the twin-modes SFG over the classical picture. Second, the number of participating mode pairs $M$ reflects the spectral bandwidth of the PDC light $\Delta\nu$. In an OPO cavity, $M$ is simply the number of cavity comb-teeth within the gain bandwidth, dictated by phase-matching. In our experiment, where the spectrum is continuous, $M\!=\!\Delta\nu T/2$ reflects the time-bandwidth product, where $\Delta\nu\!\approx\!113$THz and the effective measurement time $T\!=\!1/\delta\nu$ reflects the spectral resolution of our prism-based spectrometer $\delta\nu\!\approx\!0.5$THz, which indicates that the effective number of modes in our experiment was $M\!\approx\!60$. Clearly, the measurement time $T$ can be pushed much longer with a better spectrometer, ultimately limited only by the coherence time (linewidth) of the pump laser, which will push the number of modes to $M\!\approx\!10^{5-6}$ and beyond. And last, note that the only assumption thus far was that the signal-idler fields are weak relative to the pump, but the parametric gain was not limited, allowing $g\geq 1$, which indicates that Eq. 7 is applicable also to the high-squeezing multi-photon regime.
@@
Our goal is to calculate the number of photons $\left\langle\hat{N}_{p}\right\rangle=\left\langle\hat{a}_{p}^{\dagger}\hat{a}_{p}\right\rangle$ at the output of the second crystal, considering the effect of lock-in detection, where the phase of all modes is dithered periodically around the locking phase $\phi_{\nu}(t)\!=\!\phi_{\nu}\!+\!\phi_{d}\sin{\omega t}$ (where $\phi_{d}$ and $\omega$ are the dithering amplitude and frequency), and the SFG intensity is anti-symmetrically integrated in each period to extract $\Delta N_{p}\!=\!\int dt\left\langle\hat{N}_{p}(t)\right\rangle\sin{\omega t}$. Since the phase dithering operates only on the PDC spectrum, the pump correction $\hat{a}_{p}^{(2)}(t)$ of Eq. 9 can be broken in two: The dithered term
@@
${N}_{p}^{\mathrm{eff}}=\frac{\mathrm{i}}{\sqrt{2}}\chi t\sum_{\nu}{\left[\alpha_{r}\hat{a}_{\nu}^{(in)\dagger}\hat{a}_{-\nu}^{(in)\dagger}-\alpha_{r}^{*}\hat{a}_{\nu}^{(in)}\hat{a}_{-\nu}^{(in)}\right]=\sqrt{2}\chi t\sum_{\nu}{\text{Im}\left[\alpha_{r}^{*}\hat{a}_{\nu}^{(in)}\hat{a}_{-\nu}^{(in)}\right]}},$ (12)
@@ 5.2 Experimental Methodology
To generate an appreciable SFG signal from the entire SPDC bandwidth, it is crucial to maintain a flat spectral phase for all signal-idler frequency pairs in the ultra-broadband spectrum to ensure that every photon pair contributes constructively to the SFG process. This demands precise control of the group delay/velocity dispersion in our SU(1,1) interferometer, achieved by a prism-based shaper of the spectral-phase in the PDC path. This phase shaper, which consists of a prism pair and a reflective 4f-telescope, enables independent control of both the second and the fourth orders of dispersion (following the concept of [44]). In order to tune the dispersion to exact compensation, we measure the spectral phase in real time by observing the spectrum of the ultra-broadband SPDC at the output of the SU(1,1) interferometer with a home-built, prism-based spectrometer. The contrast of the observed SU(1,1) interference was optimized following the guidelines of [40].
@@
Finally, to isolate the desired SFG signal from the much stronger pump residue, we employ lock-in detection by dithering (modulating) the phase of the entire PDC spectrum (spontaneous and stimulated) according to $\phi(t)\!=\!\phi_{D}\sin\omega t$ using the PZT mirror in the spectral shaper (at frequency $\nu_{D}\!=\!2\pi\!\times\!173\,\mathrm{kHz}$). Due to the coherent nature of the SFG process, this phase dithering is imprinted onto the SFG result, which then interferes with the pump-residue field, modulating the detected SFG intensity, as illustrated in Fig. 5c and analyzed theoretically later on. By measuring the spectral intensity of the detected SFG signal at the modulation frequency $I(\nu)$, we effectively isolate the modulated SFG contribution from the large and noisy DC background of the pump-residue.
@@ 5.3 Theoretical Calculation of Seeded Coherent SFG
In order to quantitatively model the SFG signal with a coherent seed, as we have in the experiment, let us consider the contribution of the one seeded mode to the SFG signal in Eq. 12 separately from the other spontaneous modes. Substituting Eq. 10 into the demodulation integral $\left\langle\Delta N_{p}\right\rangle\!=\!\int dt\left\langle N_{p}(t)\right\rangle\sin{\nu t}$ we obtain the contribution of a single mode pair at frequency $\nu$
@@
The experiment measured the RF spectrum of the SFG intensity, which is proportional to $\left\langle\Delta N_{p}\right\rangle^{2}$ .
@@ 6.1 The non-classical response to loss
To gain deeper insight into the detection dynamics of our coherent SFG method and the relations between the interfering SFG contributions – spontaneous and stimulated, we investigated the response of the SFG signal to SPDC loss, regarding the theoretical prediction of Eq. 15. SPDC loss affects both photons of a pair, leading to a unique quantum situation, where although the flux of individual photons is reduced by $T$ the biphoton rate is reduced quadratically, as the probability of both photons surviving the SPDC loss is $T^{2}$.
@@
Funding This project was funded by the Israel Innovation Authority and Elta Systems Ltd. under project No. 70002 and under a Danish-IsraeliăEurekaăproject. Funding is acknowledged also from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101178170 and from the Israel Science Foundation under grant No. 2208/24.